# Few-Shot API Attack Anomaly Detection in a Classification-by-Retrieval Framework

### v1

@@ Abstract
Application Programming Interface (API) attacks refer to the unauthorized or malicious use of APIs, which are often exploited to gain access to sensitive data or manipulate online systems for illicit purposes. Identifying actors that deceitfully utilize an API poses a demanding problem. Although there have been notable advancements and contributions in the field of API security, there still remains a significant challenge when dealing with attackers who use novel approaches that don’t match the well-known payloads commonly seen in attacks. Also, attackers may exploit standard functionalities in unconventional manners and with objectives surpassing their intended boundaries. This means API security needs to be more sophisticated and dynamic than ever, with advanced computational intelligence methods, such as machine learning models that can quickly identify and respond to anomalous behavior. In response to these challenges, we propose a novel few-shot anomaly detection framework, named FT-ANN. This framework is composed of two parts: First, we train a dedicated generic language model for API based on FastText embedding. Next, we use Approximate Nearest Neighbor search in a classification-by-retrieval approach. Our framework enables the development of a lightweight model that can be trained with minimal examples per class or even a model capable of classifying multiple classes. The results show that our framework effectively improves API attack detection accuracy compared to various baselines.
Index Terms:
API Security, Anomaly Detection, Few-Shot Learning, ANN, Classification-by-Retrieval, NLP.
I Introduction
@@
Application Programming Interface (API) refers to a set of routines, procedures, resources, and protocols that permit the interaction between software systems and data exchange services  [1, 2, 3]. APIs are an evolving technology for orchestrating applications utilizing web technology [4, 5]. Recently, it has been argued that we are currently living in an API economy [5] due to the growing interconnectedness of people, applications, and systems, all of which are powered by APIs. These interfaces now serve as the foundational framework of the digital ecosystem, establishing connections between industries and economies to foster value creation and cultivate innovative capabilities [2]. APIs find utilization across a broad spectrum of services, including Web applications, Operation Systems (OS), Databases, and Hardware [6]. The increasing ease of building web applications has led to a rise in agile development, where even inexperienced engineers can deploy applications. However, this approach often lacks strong security design or hardening planning. This can lead to vulnerabilities in application logic and inadequate consideration of security impacts. For instance, failure to properly constrain resources or access levels could lead to denial of service attacks [7]. As a result, the extensive adoption of web APIs has heightened the potential of user safety and privacy breaches, making APIs a prime target for cyber attackers [3]. In recent years, there have been several high-profile API attacks [8, 9, 10, 11], such as the Zoom video conferencing platform in 2020.
@@
To address these challenges, the Open Web Application Security Project (OWASP [^1] The Open Web Application Security Project®, https://owasp.org/) provides resources, tools, and best practices to help organizations and developers enhance the security of their web applications and protect against malicious attacks. One of the most well-known contributions is the OWASP API Top 10, which outlines the ten most critical API security risks [12].
@@
Despite the progress made in prior research on utilizing machine learning and deep learning models for protecting APIs against both known and unknown attacks [13, 14, 15], there are lingering concerns that remain unresolved. These concerns encompass several challenges, such as effectively detecting zero-day vulnerabilities, minimizing false positives, and addressing real-time and continuous protection requirements. Given that a zero-day API attack involves an unknown vulnerability that the security solutions, such as web application firewalls, are unaware of, it becomes necessary to employ few-shot learning techniques. The anomaly detection model needs to leverage its comprehension of previously encountered samples to make predictions on new, unseen samples. In light of the requirement to improve traditional methods for training binary anomaly classifiers, the utilization of Classification-by-Retrieval presents a solution [16]. This approach enables the construction of neural network-based classifiers without the need for computationally intensive training procedures. Consequently, it facilitates the development of a lightweight model that can be trained with minimal examples per class or even a model capable of classifying multiple classes [17, 18, 19].
The remaining sections of this paper are structured in the following manner. In Section II, we summarize the contributions made in this paper, Section III provides an overview of the relevant literature. The architecture of the FT-ANN solution is presented in Section IV. The experimental design is outlined in Section V. Section VI focuses on measuring the effectiveness of FT-ANN and presenting experimental results in comparison to various state-of-the-art benchmarks. Finally, we draw our conclusions, current research limitations, and suggestions for future work in Section VII.
II Our Contribution
In this paper, we introduce a novel unsupervised few-shot anomaly detection framework, utilizing FastText embedding and Approximate Nearest Neighbor search (FT-ANN), which leverages a Classification-by-retrieval approach. We enable training of a single retrieval model capable of handling multiple baselines simultaneously. This approach not only saves computational resources but also simplifies the model deployment and maintenance processes. By reducing the number of models required, our methodology offers a more efficient and scalable solution for anomaly detection. Moreover, unlike traditional anomaly detection models, ANN supports incremental index updates.
@@ II Our Contribution
Additionally, we define a novel tokenizer that specifically emphasizes the language factors present in APIs, addressing the unique challenges associated with API-based natural language processing. Unlike existing tokenizers, our approach takes into account the specific language characteristics of APIs, enabling more accurate and efficient processing of API-related text. APIs are defined by a URL-based syntax in which each URL corresponds to a particular resource or action. They also include fundamental actions such as GET and POST, which determine how requests and responses are structured. Additionally, APIs utilize standard structure of HTTP headers to transmit metadata pertaining to both the request and the anticipated response.
@@
Furthermore, our language model is designed to be domain-agnostic, eliminating the need for retraining when transitioning to different API domains. This flexibility allows our model to seamlessly serve various domains without sacrificing performance or requiring additional training efforts.
@@
Lastly, its agnostic nature allows it to seamlessly adapt and address the unique requirements and challenges as also outlined in the OWASP Top 10 API vulnerabilities, posed by different API forms include REST, GraphQL, gRPC, and WebSockets, regardless of the emphasis on HTTP datasets during the demonstration.
III Related Work
Reddy et al. [20] proposed supervised sequence models based on Recurrent Neural Networks (RNN) to identify malicious injections in API requests in addition to a heuristic rule which classified 10% of each request sequence with a probability of 60% as valid to minimize the number of false positives. However, such rules can introduce subjective biases and limitations. They generated a custom-labeled dataset, but used only the request payloads. This presents a significant constraint in real-world scenarios, where malicious actor could potentially manipulate a majority of the request components, including headers [21]. They compared six unidirectional and bidirectional RNN models by evaluating various performance measures and showed a 50% decrease in false positive cases.
Jemal et al. [22] suggested the Convolutional Neural Network (CNN) method to detect web attacks. They concluded that appropriately adjusting hyper-parameters and employing a data pre-processing approach significantly impacts the detection rate.
Gniewkowski et al. [23] suggested an NLP-based semi-supervised anomaly detection methodology that employs RoBERTa for embedding HTTP requests to detect anomalies. They evaluated the pipeline over the CSIC 2010 and CSE-CIC-IDS 2018 published datasets, in addition to UMP, a custom-generated dataset, achieving F1-scores of 96.9%, 92.6%, and 99.9%, respectively. A notable limitation in their method is the relatively long training and inference time, which presents a considerable drawback in real-time systems.
@@ III Related Work
Jemal et al. [24] propose a supervised Memory CNN that combines a CNN and Long Short-Term Memory (LSTM) to identify patterns of malicious requests within sequences of requests. They evaluated the model using the CSIC 2010 dataset. Niu et al. [25] propose a technique for detecting web attacks based on a supervised CNN and Gated Recurrent Unit (GRU). They extracted statistical features and utilized a Word2Vec model to extract word embeddings, resulting in a 3-dimensional input for the suggested CNN-GRU method. A fully connected layer was used for classification, and the experiment was made on the CSIC 2010 dataset. Yu et al. [26] combined a CNN and Support Vector Machine (SVM) to detect malicious web server requests. The CSIC 2010 dataset was selected to validate proposed approach. Baye et al. [14] also utilized SVM with a Linear Kernel as a two-class classifier to identify anomalous API requests. To form a training dataset that accurately represented authentic API logs, they employed a technique for outlier detection based on Gaussian Distribution, generating a synthetic dataset with labeled examples.
Moradi et al.[27] introduced an unsupervised anomaly detection technique utilizing Auto-Encoder LSTM for feature extraction and Isolation Forest for classification. They applied this model to the CSIC 2010 dataset, achieving an F1-score of 81.96%. Although, they faced limitations related to the choice of encoding method for HTTP data, the non-stationary nature of HTTP data, and the integration of multiple feature sets, which may impact the effectiveness of the proposed approach. In more recent research [28], they also proposed an unsupervised Deep Support Vector method. They conducted a comparison between two feature extraction approaches, namely bigram (2-gram) and one-hot. They evaluated it on the CSIC 2010 and ECML/PKDD 2007 datasets, achieving F1-scores of 89% and 79.48%, respectively. They pointed out the limitation of lacking support for data streams in incremental learning, emphasizing the need for future research to address this aspect.
Ref Attack Method Task Technique(s) Method(s) Dataset(s) Best Score
 [20] SQL, XML and JSON attacks in HTTP request Supervised RNN Word embedding Custom 98.13% F1-Score
 [22] Attack in HTTP request Supervised CNN Word embedding, Character embedding CSIC 2010 97.65% Accuracy
 [23] Attack in HTTP request Semi-Supervised RoBERTa BBPE CSIC 2010 96.9% F1-Score
UMP 92.6% F1-Score
CSE-CIC-IDS2018 99.9% F1-Score
 [24] Sequences of HTTP requests attack Supervised CNN, LSTM ASCII embedding CSIC 2010 98.53% F1-Score
 [25] Attack in HTTP request Supervised CNN, GRU Word embedding, Statistical features CSIC 2010 98.77% F1-Score
 [26] Attack in HTTP request Supervised TextCNN, SVM Word embedding,t-SNE, Statistical features CSIC 2010 99.3% F1-Score
 [27] Attack in HTTP request Unsupervised Auto-Encoder LSTM Character embedding CSIC 2010 81.96% F1-Score
 [28] Attack in HTTP request Unsupervised SVDD Character embedding CSIC 2010 89% F1-Score
@@
TABLE I: Summary of Machine Learning NLP-based Techniques
IV Framework
@@
The analysis of API requests can be framed as a problem in NLP. One challenge lies in selecting a language model capable of generating a vector space representation. In our work, we decided to utilize the FastText [29] model. The primary aim of FastText embeddings is to factor in the internal structure of words rather than simply learning word representations. This feature is especially advantageous for morphologically complex languages, allowing representations for various morphological forms of words to be learned separately. FastText offers Skip-gram and Continuous Bag-of-Words (CBOW) models to compute word representations. Although CBOW learns faster than Skip-gram, Skip-gram outperforms CBOW on small datasets [30]. FastText adopts a character n-gram approach to tokenize words, which effectively tackles Out-of-vocabulary (OOV) problems. This method not only generates embeddings for common words but also for rare, misspelled, or previously unseen words in the training corpus. In the realm of API security, it is imperative to pay attention to the internal structure of words in order to grasp the intent and context behind API requests. API security entails scrutinizing the textual content within API requests to pinpoint potential threats or vulnerabilities [31]. These security threats could be concealed within apparently harmless text, which may involve the use of uncommon or previously unseen terms, or even attempts to obscure their intent through spelling errors [32]. Moreover, APIs might be required to accommodate a diverse range of languages, including those characterized by intricate morphological structures [33].
@@
Prior works have proposed various methods for detecting anomalies. While the majority of these approaches focus on classic classification tasks, we propose the utilization of the ANN vector similarity method for identifying in-distribution records. The similarity search concept pertains to the process of identifying data points in a dataset that demonstrate similarities with a specific pattern, commonly referred to as a query [34]. To measure the similarity of a pair of data points, a distance function is used, where a small distance indicates that the two points are more similar or ”closer” to each other [35]. The NN search is a specific type of similarity search used to identify data points that are nearest in distance to a provided query point [36]. The ANN allows search despite the possibility of not retrieving all neighbors in a metric space [36].
@@
While both ANN and traditional NN are rooted in the fundamental concept of similarity search, traditional NN search involves an exhaustive examination of all data points to find the NN, which can be prohibitively time-consuming for large datasets [36]. In contrast, ANN employs techniques that trade off a slight loss in precision for substantial gains in speed, enabling the identification of ANN without examining every data point. This efficiency becomes especially crucial in real-world API applications where rapid response times are essential. Additionally, ANN methods are adaptable to high-dimensional spaces [34, 37], where traditional NN searches can suffer from the ”curse of dimensionality”, in which the computational requirements for exact NN search become prohibitively high as the dimensionality of the dataset increases [36].
@@
The framework proposed (depicted in Figure 1) comprises a combination of a FastText embedding network and a retrieval layer, which includes an ANN matching component and a result aggregation component built upon it, forming the FT-ANN system. In Phase 1 (Figure 1), we initiate the process by training a generic language model to serve as a reliable baseline for the detection model. To gather a substantial dataset, we conducted web crawling on a random sample of websites from the Tranco top websites list, collecting over a million examples of normal web traffic. This language model now serves as a pre-trained baseline, applicable to any real-world API anomaly detection system.
@@
Proceeding to Phase 2 (Figure 1), the unsupervised detection model is trained, consisting of a pre-processing step and a Classification-by-retrieval framework. The pre-processing step validates the HTTP headers against the standard structure. Additionally, a unique data transformation is applied to simplify the API vocabulary. The detection model is trained solely on normal API traffic, and constructs a single ANN model for all endpoints. The term ”endpoint” refers to any data or metadata that may represent a type, an origin, or identification of a respective API request. Endpoint is defined as the combination of the method, host and path [38]. For each endpoint, a threshold is calculated and utilized during the detection stage.
@@
Finally, in Phase 3 (Figure 1), the model’s performance is validated by allowing an index search for every request. During validation, we standardize the request using the same pre-processing stages applied during the detection model training phase. The text is then transformed into a vector representation using our generic language model through inference. These embedding vectors are used in the ANN search, which returns the K-NNs from the specific endpoint index collection, followed by a maximum distance scaling layer. The final score obtained enables the model to ascertain whether the incoming API request is normal or an anomaly by comparing with the pre-defined threshold for each API endpoint.
Figure 1: Our FT-ANN Framework
IV-A Data Pre-Processing

### v2

@@ Abstract
Application Programming Interface (API) Injection attacks refer to the unauthorized or malicious use of APIs, which are often exploited to gain access to sensitive data or manipulate online systems for illicit purposes. Identifying actors that deceitfully utilize an API poses a demanding problem. Although there have been notable advancements and contributions in the field of API security, there remains a significant challenge when dealing with attackers who use novel approaches that don’t match the well-known payloads commonly seen in attacks. Also, attackers may exploit standard functionalities unconventionally and with objectives surpassing their intended boundaries. Thus, API security needs to be more sophisticated and dynamic than ever, with advanced computational intelligence methods, such as machine learning models that can quickly identify and respond to abnormal behavior. In response to these challenges, we propose a novel unsupervised few-shot anomaly detection framework composed of two main parts: First, we train a dedicated generic language model for API based on FastText embedding. Next, we use Approximate Nearest Neighbor search in a classification-by-retrieval approach. Our framework allows for training a fast, lightweight classification model using only a few examples of normal API requests. We evaluated the performance of our framework using the CSIC 2010 and ATRDF 2023 datasets. The results demonstrate that our framework improves API attack detection accuracy compared to the state-of-the-art (SOTA) unsupervised anomaly detection baselines.
keywords:
[…]
{highlights}
A novel unsupervised few-shot anomaly detection framework to detect API Injection attacks.
Introduced a tokenizer designed to capture and emphasize language factors specific to APIs, addressing unique NLP challenges.
Leveraged FastText embedding combined with Approximate Nearest Neighbor search, employing a Classification-by-retrieval approach.
Validated the framework using public HTTP datasets against cutting-edge techniques.
1 Introduction
@@
Application Programming Interface (API) refers to a set of routines, procedures, resources, and protocols that permit the interaction between software systems and data exchange services  (Balsari et al., 2018; Ofoeda et al., 2019; Mendoza and Gu, 2018). APIs are an evolving technology for orchestrating applications utilizing web technology (Benzaid and Taleb, 2020; IBM, 2016). Recently, it has been argued that we live in an API economy (IBM, 2016) due to the growing interconnectedness of people, applications, and systems, all of which are powered by APIs. These interfaces now serve as the foundational framework of the digital ecosystem, establishing connections between industries and economies to foster value creation and cultivate innovative capabilities (Ofoeda et al., 2019). APIs find utilization across a broad spectrum of services, including Web applications, Operation Systems (OS), Databases, and Hardware (Reddy, 2011). The increasing ease of building web applications has led to a rise in agile development, where even inexperienced engineers can deploy applications. However, this approach often lacks strong security design or hardening planning. This can lead to vulnerabilities in application logic and inadequate consideration of security impacts. For instance, failure to properly constrain resources or access levels could lead to denial of service attacks (Sun et al., ). As a result, the extensive adoption of web APIs has heightened the potential of user safety and privacy breaches, making APIs a prime target for cyber attackers (Mendoza and Gu, 2018). In recent years, there have been several high-profile API attacks (Coyne, 2020; BBC, 2019; Greig, 2023; Tara Siegel et al., 2017), such as the Zoom video conferencing platform in 2020.
@@
To address these challenges, the Open Web Application Security Project (OWASP [^1] https://owasp.org/) provides resources, tools, and best practices to help organizations and developers enhance the security of their web applications and protect against malicious attacks. One of the most well-known contributions is the OWASP API Top 10, which outlines the ten most critical API security risks (Fett et al., 2016). Injection flaws enable attackers to inject malicious code through an application into another system. These flaws encompass a variety of attacks, such as Directory Traversal, SQL Injection (SQLi), LDAP Injection, and Cross-Site Scripting (XSS). OWASP classifies Injection flaws under the broader category of Unsafe Consumption of APIs security risks [^2] https://owasp.org/API-Security/editions/2023/en/0xaa-unsafe-consumption-of-apis/.
@@
Despite the progress made in prior research on utilizing machine learning and deep learning models for protecting APIs against both known and unknown attacks (Chan et al., 2023; Baye et al., 2021; Harlicaj et al., 2021), there are lingering concerns that remain unresolved. These concerns encompass several challenges, such as effectively detecting zero-day vulnerabilities, minimizing false positives, and addressing real-time and continuous protection requirements. Given that a zero-day API attack involves an unknown vulnerability that the security solutions, such as web application firewalls, are unaware of, it becomes necessary to employ few-shot learning techniques. The anomaly detection model needs to leverage its comprehension of previously encountered samples to make predictions on new, unseen samples. In light of the requirement to improve traditional methods for training binary anomaly classifiers, the utilization of Classification-by-Retrieval presents a solution (Shen et al., 2017). This approach enables the construction of neural network-based classifiers without requiring computationally intensive training procedures. Consequently, it facilitates the development of a lightweight model that can be trained with minimal examples per class or even a model capable of classifying multiple classes (Shi et al., 2022; Qamar et al., 2008; Valero-Mas et al., 2023). In this context, we propose a novel unsupervised few-shot anomaly detection framework, utilizing FastText embedding and Approximate Nearest Neighbor search (FT-ANN). This framework offers a unique representation for the API request, facilitating intelligent segmentation based on the API endpoint. The term "endpoint" refers to any data or metadata that may represent a type, an origin, or identification of a respective API request. Endpoint is defined as the combination of the method, host, and path (Battle and Benson, 2008). This paper’s contributions are itemized as follows:
  1. [•]
  2. 1.
    We introduce a novel unsupervised few-shot anomaly detection framework, utilizing FastText embedding and ANN search, which leverages a Classification-by-retrieval approach. This framework employs a Classification-by-retrieval approach and allows for training a single retrieval model that can handle multiple baselines concurrently, such as API endpoints or multiple API domains. Our approach provides a more efficient and scalable solution for anomaly detection by reducing the number of required models and supporting incremental index updates. This is in contrast to traditional models, which usually require the entire dataset to be present for training new examples. Furthermore, the high performance and speed of the in-memory similarity search, widely adopted in many ecosystems due to its flexible nature, significantly enhance the overall efficiency of our framework.
  3. 2.
@@
    We define a novel tokenizer that specifically emphasizes the language factors present in APIs, addressing the unique challenges associated with API-based natural language processing. APIs are defined by a URL-based syntax in which each URL corresponds to a particular resource or action, and they include fundamental actions such as GET and POST, determining the structure of requests and responses. Additionally, APIs utilize the standard structure of HTTP headers to transmit metadata pertaining to both the request and the response. Unlike existing tokenizers, our approach considers these specific language characteristics of APIs, enabling more accurate and efficient processing of API-related text.
  4. 3.
@@
    Our language model is designed to be domain-agnostic, eliminating the need for retraining when transitioning to different API domains. This flexibility allows our model to seamlessly serve various domains without sacrificing performance or requiring additional training efforts.
  5. 4.
@@
    Our model’s agnostic nature allows it to seamlessly adapt and address the unique requirements and challenges, as also outlined in the OWASP Top 10 API vulnerabilities, posed by different API forms, including REST, GraphQL, gRPC, and WebSockets, regardless of the emphasis on HTTP datasets during the demonstration.
The remaining sections of this paper are structured in the following manner. In Section 2 we provide an overview of the relevant literature; The architecture of the FT-ANN framework is outlined in Section 3; Section 4 describes the datasets that were used for this paper; The experimental design is outlined in Section 5; Section 6 outlines the experimental results, which focus on measuring the effectiveness the FT-ANN framework and presenting experimental results compared to various state-of-the-art benchmarks; Finally, we draw conclusions, current research limitations, and suggestions for future work in Section 7. For ease of reading, we provide a list of common abbreviations in Table 1.
Abbreviation Definition
AE Autoencoder
ANN Approximate Nearest Neighbor
API Application Programming Interface
CBLOF Clustering-Based Local Outlier Factor
CNN Convolutional Neural Network
DeepSVDD Deep Support Vector Data Description
FB Feature Bagging
GRU Gated Recurrent Unit
HBOS Histogram-based Outlier Detection
IF Isolation Forest
KDE Kernel Density Estimation
LMDD Linear Model Deviation-based Outlier Detection
LOF Local Outlier Factor
LSTM Long Short-Term Memory
MCD Minimum Covariance Determinant
NLP Natural Language Processing
OCSVM One-class SVM
PCA Principal Component Analysis
QPS Queries per second
RDA Regularized Deep Autoencoder
RNN Recurrent Neural Networks
SOTA State-of-the-art
SQLi SQL Injection
SVM Support Vector Machine
XSS Cross-Site Scripting
Table 1: List of Abbreviations.
2 Related Work
This section reviews relevant research on API and web traffic classification, focusing on the use of machine learning, deep learning, and natural language processing (NLP) techniques.
@@ 2 Related Work
Reddy and Rudra (2021) proposed supervised sequence models based on Recurrent Neural Networks (RNN) to identify malicious injections in API requests in addition to a heuristic rule that classified 10% of each request sequence with a probability of 60% as valid to minimize the number of false positives. However, such rules can introduce subjective biases and limitations. They generated a custom-labeled dataset but used only the request payloads. This presents a significant constraint in real-world scenarios, where malicious actors could potentially manipulate a majority of the request components, including headers (Ombagi, 2017). They compared six unidirectional and bidirectional RNN models by evaluating various performance measures and showed a 50% decrease in false positive cases. Jemal et al. (2021a) suggested the Convolutional Neural Network (CNN) method to detect web attacks. They concluded that appropriately adjusting hyper-parameters and employing a data pre-processing approach significantly impacts the detection rate. Jemal et al. (2020) propose a supervised Memory CNN that combines a CNN and Long Short-Term Memory (LSTM) to identify patterns of malicious requests within sequences of requests. Niu and Li (2020) propose a technique for detecting web attacks based on a supervised CNN and Gated Recurrent Unit (GRU). They extracted statistical features and utilized a Word2Vec model to extract word embeddings, resulting in a 3-dimensional input for the suggested CNN-GRU method. A fully connected layer was used for classification. Yu et al. (2020) combined a CNN and Support Vector Machine (SVM) to detect malicious web server requests. Baye et al. (2021) also utilized SVM with a Linear Kernel as a two-class classifier to identify anomalous API requests. To form a training dataset that accurately represented authentic API logs, they employed a technique for outlier detection based on Gaussian Distribution, generating a synthetic dataset with labeled examples.
Gniewkowski et al. (2021) suggested an NLP-based self-supervised anomaly detection methodology that employs RoBERTa for embedding HTTP requests. A notable limitation in their method is the relatively long training and inference time, which presents a considerable drawback in real-time systems.  Moradi Vartouni et al. (2019) introduced an unsupervised anomaly detection technique utilizing Autoencoder (AE) and LSTM for feature extraction and Isolation Forest (IF) for classification. However, they faced limitations related to the choice of encoding method for HTTP data, the non-stationary nature of HTTP data, and the integration of multiple feature sets, which may impact the effectiveness of the proposed approach. In a more recent study (Vartouni et al., 2023), an unsupervised Deep Support Vector Data Description (DeepSVDD) method was proposed. They conducted a comparison between two feature extraction approaches, namely bigram (2-gram) and one-hot. They pointed out the limitation of lacking support for data streams in incremental learning, emphasizing the need for future research to address this aspect. Mac et al. (2018) also suggested employing AEs for unsupervised learning, evaluated various variants, and concluded that the Regularized Deep AE (RDA) performed the best. They processed the raw HTTP requests and tokenized the URLs by replacing the characters with their corresponding ASCII codes. Vartouni et al. (2018) used a Stacked AE to extract HTTP features and employed IF as the classifier. In a Stacked AE, the output of each AE serves as the input for the next layer. They demonstrated that using various activation functions and optimization techniques can improve the learning process. Demirel and Sandikkaya (2023) proposed combining the Local Outlier Factor (LOF) algorithm and an AE to enhance classification accuracy. The study evaluates different combination approaches, including the Behavior Knowledge Space (BKS) for combining multiple classifiers based on their predictions and Weighted Borda Counts (wBorda) for weighted voting systems, to assess their effectiveness.
Ref Task Technique(s) Method(s) Dataset(s)
 Yu et al. (2020) Supervised TextCNN, SVM Word embedding,t-SNE, Statistical features CSIC 2010
 Niu and Li (2020) Supervised CNN, GRU Word embedding, Statistical features CSIC 2010
 Jemal et al. (2020) Supervised CNN, LSTM ASCII embedding CSIC 2010
 Reddy and Rudra (2021) Supervised RNN Word embedding Custom
 Jemal et al. (2021a) Supervised CNN Word embedding, Character embedding CSIC 2010
 Gniewkowski et al. (2021) Self-supervised RoBERTa BBPE, Doc2vec CSIC 2010
Custom
CSE-CIC-IDS2018
 Mac et al. (2018) Unsupervised AE ASCII embedding CSIC 2010
 Vartouni et al. (2018) Unsupervised SAE, IF Character N-grams CSIC 2010
 Moradi Vartouni et al. (2019) Unsupervised AE LSTM Character embedding CSIC 2010
 Vartouni et al. (2023) Unsupervised DeepSVDD Character embedding CSIC 2010
ECML/PKDD 2007
 Demirel and Sandikkaya (2023) Unsupervised AE, LOF ASCII embedding CSIC 2010
Custom
Fwaf
FT-ANN Unsupervised Classification-by-retrieval Sentence embedding CSIC 2010
ATRDF 2023
@@
Table 2: Overview of machine learning and NLP-based techniques for detecting attacks in HTTP traffic.
3 Framework
@@
The analysis of API requests can be framed as a problem in NLP. One challenge lies in selecting a language model capable of generating a vector space representation. In our work, we decided to utilize the FastText (Research, 2017) model. FastText embeddings primarily aim to factor in the internal structure of words rather than simply learning word representations. This feature is especially advantageous for morphologically complex languages, allowing representations for various morphological forms of words to be learned separately. FastText offers Skip-gram and Continuous Bag-of-Words (CBOW) models to compute word representations. Although CBOW learns faster than Skip-gram, Skip-gram outperforms CBOW on small datasets (Bansal and Srivastava, 2018). FastText adopts a character n-gram approach to tokenize words, which effectively tackles Out-of-vocabulary (OOV) problems. This method not only generates embeddings for common words but also for rare, misspelled, or previously unseen words in the training corpus. In the realm of API security, it is essential to examine the internal structure of requests to comprehend their purpose and context. Security threats might be hidden within seemingly innocuous text, which could include the use of uncommon or previously unseen terms, or even efforts to mask their intention through spelling errors (Xiao et al., 2016). Moreover, APIs might be required to accommodate a diverse range of languages, including those characterized by intricate morphological structures (Seyyar et al., 2022).
@@
Prior works have proposed various methods for detecting anomalies. While most of these approaches focus on classic classification tasks, we propose utilizing the ANN vector similarity method for identifying in-distribution records. The similarity search concept pertains to the process of identifying data points in a dataset that demonstrate similarities with a specific pattern, commonly referred to as a query (Chávez et al., 2001). To measure the similarity of a pair of data points, a distance function is used, where a small distance indicates that the two points are more similar or "closer" to each other (Ponomarenko et al., 2014). The Nearest Neighbor (NN) search is a specific type of similarity search used to identify data points nearest in distance to a provided query point (Indyk and Motwani, 1998). The ANN allows search despite the possibility of not retrieving all neighbors in a metric space (Indyk and Motwani, 1998).
@@
While both ANN and traditional NN are rooted in the fundamental concept of similarity search, traditional NN search involves an exhaustive examination of all data points to find the NN, which can be prohibitively time-consuming for large datasets (Indyk and Motwani, 1998). In contrast, ANN employs techniques that trade-off a slight loss in precision for substantial gains in speed, enabling the identification of ANN without examining every data point. This efficiency becomes especially crucial in real-world API applications where rapid response times are essential. Additionally, ANN methods are adaptable to high-dimensional spaces (Chávez et al., 2001; Hajebi et al., 2011), where traditional NN searches can suffer from the "curse of dimensionality", in which the computational requirements for exact NN search become prohibitively high as the dimensionality of the dataset increases (Indyk and Motwani, 1998).
@@
The framework proposed (depicted in Figure 1) comprises a combination of a FastText embedding network and a retrieval layer, which includes an ANN matching component and a result aggregation component built upon it, forming the FT-ANN system. In Phase 1 (Figure 1), we initiate the process by training a generic language model to serve as a reliable baseline for the detection model. To gather a substantial dataset, we crawled a random sample of websites from the Tranco top websites list, collecting over a million examples of normal web traffic. This language model now serves as a pre-trained baseline, applicable to any real-world API anomaly detection system.
@@
Proceeding to Phase 2 (Figure 1), the unsupervised detection model is trained, consisting of a pre-processing step and a Classification-by-retrieval framework. The pre-processing step validates the HTTP headers against the standard structure. Additionally, a unique data transformation is applied to simplify the API vocabulary. The detection model is trained solely on normal API traffic and constructs a single ANN model for all endpoints. For each endpoint, a threshold is calculated and utilized during the detection stage.
@@
Finally, in Phase 3 (Figure 1), the model’s performance is validated by allowing an index search for every request. During validation, we standardize the request using the same pre-processing stages applied during the detection model training phase. The text is then transformed into a vector representation using our generic language model through inference. These embedding vectors are used in the ANN search, which returns the K-NNs from the specific endpoint index collection, followed by a maximum distance scaling layer. The final score obtained enables the model to ascertain whether the incoming API request is normal or an anomaly by comparing it with the pre-defined threshold for each API endpoint.
Figure 1: The FT-ANN Framework: Training and Inference Methodology.
3.1 Data Pre-Processing